the sum of 2 numbers is 55.the difference is 7. what are the numbers?
Answer by TimothyLamb(4379)

n + m = 55
m - n = 7
m = 7 + n
---
n + m = 55
n + 7 + n = 55
2n = 55 - 7
2n = 48
---
n = 24
m = 31
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php